If the car passes point A with a speed of 20m/s and begins to increase its speed at a constant rate of at=0.5m/s2, determine the magnitude of the car’s acceleration when S = 101.68 m and x = 0.

730m—
Answer #1

PAGE NO DATE To calculate the magnitude of the Car at S= 101.68 m. I V = V + 295 - put the values. Vaê 2omis a = 0.5 m/s2 S=

Radius of curvatune 8 = [1 + (-2x 1625) 273/2 - 2 625 At x = 0 8 = [lto]3/2 + 0 -625 = - 312. 5m 312.5m anz o Vs g (Normal co

Tangential component = 0.5 mls Acceleration az Jaetan - 10.5²+ lobos ² a = 11:69 m/s2